4x+y-z=8
x-y+2z=3
3x-y+z=6
I am not even sure where to start with this problem....

Answer by checkley77(12844)

4x+y-z=8 add equations 1 & 3
x-y+2z=3
3x-y+z=6
4x+y-z=8
----------
7x=14
x=14/2
x=2 ans.
Replace x with 2 in equation 2 & 3
2-y+2z=3
-y+2z=3-2
-y+2z=1 subtract this equation from the last equation.
3*2-y+z=6
6-y+z=6
-y+z=6-6
-y+z=0
-(-y+2z=1)
--------------
-z=-1
z=1 ans.
4*2+y-1=8
8+y-1=8
y=8-8+1
y=1 ans.
x,y,z (2,1,1)